# Quadrupole-hexadecapole correlations in neutron-rich samarium and gadolinium isotopes

### v1

@@ IV.2.3 Monopole transitions
One can see from Figs. 12(a) and (c) that for those Sm and Gd nuclei with $N\geq 92$, the $sdg$-IBM predicts the monopole strengths $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ to be more or less similar to eath other, whereas in the $sd$-IBM these monopole transition strengths are quite at variance, that is, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ are larger than the $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ by a factor of 2 to 3. In addition, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ values obtained from the mapped $sdg$-IBM are smaller than those from the $sd$-IBM for neutron-rich nuclei with $N\geq 94$. The reduction of the $0_{2}^{+}\rightarrow 0_{1}^{+}$ $E0$ transition in the case of the $sdg$-IBM has been already been discussed in Ref. [44]. On the other hand, the $sdg$-IBM provides larger $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ values for the $N\geq 94$ nuclei than the $sd$-IBM. These quantitative differences in the predicted $\rho^{2}(E0)$ values between the two sets of the IBM calculations represent the most significant difference between the $sdg$-IBM and $sd$-IBM in the neutron-rich Sm and Gd isotopes with $N\geq 94$. As in the case of the hexadecapole transitions, experimental information about the monopole transitions is limited. It is therefore not possible to tell whether the $sdg$-IBM represents an improvement over the simpler $sd$-IBM in this respect.
@@ V Summary
The mapped $sdg$-IBM improves the description of the high-spin yrast states in the nuclei with $N=88$, as well as in the Sm isotopes with $N=90$ and 92. In heavier rare-earth nuclei, the contribution of $g$ bosons to the ground state band is insignificant. In the case of the excited $K=0^{+}$ and $\gamma$-vibrational bands, the $sdg$-IBM results did not differ significantly from the $sd$-IBM ones, with only some minor improvements of the $4_{3}^{+}$ states of the $K=0^{+}$ band at $N=88$ in the $sdg$-IBM. In well-deformed $N=90$ and 92 nuclei, the $sdg$-IBM predicts stronger $B(E2)$ transition strengths for the yrast states with spin $J^{\pi}\geq 6^{+}$, which is in good agreement with the experimental data. The effect is more visible in 152,154Sm, for which pronounced hexadecapole mean-field minima were found in the PESs. The choice of the Gogny D1S EDF seems to lead to somewhat stronger yrast band $B(E2)$ values compared to the ones obtained with the relativistic DD-PC1 EDF. The existence of the $K=4^{+}$ band with an enhanced $B(E4;4^{+}\rightarrow 0^{+})$ transition to the ground state is predicted by the $sdg$-IBM in 150Sm and 152,154Gd. In heavier isotopes, such bands are also predicted, however, the bandhead in those nuclei is predicted to be a higher-lying $4_{n\geq 5}^{+}$ state. Regarding monopole strengths, the $sdg$-IBM predicts lower values of $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and larger values of $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ in neutron-rich isotopes with $N\geq 94$, than the $sd$-IBM. This is in good agreement with previous theoretical calculations, and represents the most pronounced hexadecapole correlation effect on the low-lying spectra in very neutron-rich Sm and Gd isotopes. Due to the lack of experimental data on such transitions, it remains to be seen whether the $sdg$-IBM correctly predicts the behaviour of monopole transitions in those isotopes.
@@ Acknowledgements.
Most of the work by L.L. was done at the University of Zagreb and was funded within the Tenure Track Pilot Programme of the Croatian Science Foundation and the École Polytechnique Fédérale de Lausanne, and the project TTP-2018-07-3554 Exotic Nuclear Structure and Dynamics, with funds from the Croatian-Swiss Research Programme. This work has also been supported by the Spanish Agencia Estatal de Investigación (AEI) of the Ministry of Science and Innovation (MCIN) under grant agreement No. PID2021-127890NB-I00. L.L. acknowledges support by the “Ramón y Cajal” grant No. RYC2021-031880-I funded by MCIN/AEI/10.13039/501100011033 and the European Union-”NextGenerationEU”. The work of R. Rodríguez-Guzmán was partially supported through the grant PID2022-136228NB-C22 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE and ”ERDF A way of making Europe”.

### v2

@@ IV.2.3 Monopole transitions
One can see from Figs. 12(a) and (c) that for those Sm and Gd nuclei with $N\geq 92$, the $sdg$-IBM predicts the monopole strengths $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ to be more or less similar to each other, whereas in the $sd$-IBM these monopole transition strengths are quite at variance, that is, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ are larger than the $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ by a factor of 2 to 3. In addition, the $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ values obtained from the mapped $sdg$-IBM are smaller than those from the $sd$-IBM for neutron-rich nuclei with $N\geq 94$. The reduction of the $0_{2}^{+}\rightarrow 0_{1}^{+}$ $E0$ transition in the case of the $sdg$-IBM has been already been discussed in Ref. [44]. On the other hand, the $sdg$-IBM provides larger $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ values for the $N\geq 94$ nuclei than the $sd$-IBM. These quantitative differences in the predicted $\rho^{2}(E0)$ values between the two sets of the IBM calculations represent the most significant difference between the $sdg$-IBM and $sd$-IBM in the neutron-rich Sm and Gd isotopes with $N\geq 94$. As in the case of the hexadecapole transitions, experimental information about the monopole transitions is limited. It is therefore not possible to tell whether the $sdg$-IBM represents an improvement over the simpler $sd$-IBM in this respect.
@@ V Summary
The mapped $sdg$-IBM improves the description of the high-spin yrast states in the nuclei with $N=88$, as well as in the Sm isotopes with $N=90$ and 92. In heavier rare-earth nuclei, the contribution of $g$ bosons to the ground state band is insignificant. In the case of the excited $K=0^{+}$ and $\gamma$-vibrational bands, the $sdg$-IBM results did not differ significantly from the $sd$-IBM ones, with only some minor improvements of the $4_{3}^{+}$ states of the $K=0^{+}$ band at $N=88$ in the $sdg$-IBM. In well-deformed $N=90$ and 92 nuclei, the $sdg$-IBM predicts stronger $B(E2)$ transition strengths for the yrast states with spin $J^{\pi}\geq 6^{+}$, which is in good agreement with the experimental data. The effect is more visible in 152,154Sm, for which pronounced hexadecapole mean-field minima were found in the PESs. The choice of the Gogny D1S EDF seems to lead to somewhat stronger yrast band $B(E2)$ values compared to the ones obtained with the relativistic DD-PC1 EDF. The existence of the $K=4^{+}$ band with an enhanced $B(E4;4^{+}\rightarrow 0^{+})$ transition to the ground state is predicted by the $sdg$-IBM in 150Sm and 152,154Gd. In heavier isotopes, such bands are also predicted; however, the bandhead in those nuclei is predicted to be a higher-lying $4_{n\geq 5}^{+}$ state. Regarding monopole strengths, the $sdg$-IBM predicts lower values of $\rho^{2}(E0;0_{2}^{+}\rightarrow 0_{1}^{+})$ and larger values of $\rho^{2}(E0;0_{3}^{+}\rightarrow 0_{1}^{+})$ in neutron-rich isotopes with $N\geq 94$ than the $sd$-IBM. This is in good agreement with previous theoretical calculations, and represents the most pronounced hexadecapole correlation effect on the low-lying spectra in very neutron-rich Sm and Gd isotopes. Due to the lack of experimental data on such transitions, it remains to be seen whether the $sdg$-IBM correctly predicts the behavior of monopole transitions in those isotopes.
@@ Acknowledgements.
Most of the work by L.L. was done at the University of Zagreb and was funded within the Tenure Track Pilot Programme of the Croatian Science Foundation and the École Polytechnique Fédérale de Lausanne, and the project TTP-2018-07-3554 Exotic Nuclear Structure and Dynamics, with funds from the Croatian-Swiss Research Programme. This work has also been supported by the Spanish Agencia Estatal de Investigación (AEI) of the Ministry of Science and Innovation (MCIN) under Grant Agreement No. PID2021-127890NB-I00. L.L. acknowledges support by the “Ramón y Cajal” Grant No. RYC2021-031880-I funded by MCIN/AEI/10.13039/501100011033 and the European Union-”NextGenerationEU”. The work of R.R.-G. was partially supported through Grant No. PID2022-136228NB-C22 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE and ”ERDF A way of making Europe.”